OpenStudy (beccab003):

Choose the correct equation of a line, in point-slope form, given the point (-1,3) and a slope, m, of 1/2. 1) y+1=1/2(x-3) 2) y-3=1/2(x+1) 3) y+3=1/2(x-1) 4) y-1=1/2(x-3)

OpenStudy (aravindg):

Equation in point slope form is given by: \(y-y_1=m(x-x_1)\) where \[(x_1,y_1) \]is the point and m is the slope.

OpenStudy (beccab003):

So is the answer the second or third?

OpenStudy (beccab003):

it has to be y-3 or y+3 because 3 is the y coordinate in the equation.

OpenStudy (aravindg):

Why do you have a confusion? y-y1 y1=?

OpenStudy (beccab003):

So is the answer the third one?

OpenStudy (beccab003):

the 3 is positive